• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 11:53:55 on October24,2005

    Entry number 157868

    keyword=Hall in controlled access, right magnets and septum temperature.

    Hall A has been in controlled access since around 11:00 for cryo work
    on the right dipole. Ed was in the hall. Jack also went in to the hall.
    Right magnets all went to zero.
    Per Jack's request, I pushed the green reset button for the right spectrometer twice.
    Right septum temperature went to above 8 K. Current went to zero.
    Informed Ed Folts. Now the right septum temperature started to coming down,